Circular No. 6689

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cs IN ANONYMOUS GALAXY J. Mueller reports her discovery of a supernova (mag ∼ 17) on a IV-N

plate taken on June 26 UT with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrogram of SN 1997cs, taken by D. Koranyi on June 30.29 UT at the Whipple Observatory 1.5-m telescope, shows this to be a type-IIn supernova with moderately broad Balmer emission (1800 km/s FWHM), along with He I emission at 587.5 and 706.5 nm. Blended emission lines of permitted Fe II are also visible. The supernova is at redshift 0.037, based on emission lines of the host galaxy. Images taken concurrently by M. Gomez with the 1.2-m telescope yield magnitude V = 17.4. G. V. Williams provides the following precise position for the supernova from one of the Gomez images: α = 15^{h}13^{m}39^{s}.78, δ = +2^{o}53′44′′.5 (equinox 2000.0; uncertainty ± 0′′.4); the corresponding offset is 18′′.1 east and 1′′.2 north of the galaxy’s center.

MXB 1730–335 R. Guerriero, W. H. G. Lewin, and J. Kommers, Massachusetts In-

stitute of Technology, report: “Following detection of the rapid burster MXB 1730–335 with the RXTE’s All-Sky Monitor, we observed the source on June 26.19 and 26.33 UT for 1500 and 2300 s, respectively, with the RXTE PCA. A persistent emission level of 4100 ± 300 counts/s (total PCA countrate, 2–60 keV; 1 Crab = 13 500 counts/s) was inferred for the rapid burster, after subtracting a contribution of 1200 ± 200 counts/s from 4U 1728–34. Seven type-II bursts were detected, with peak countrates of 1100–2900 counts/s above the persistent emission. The bursts had a 10–20s smooth rise and a 40–60-s decay, and occurred 450–650 s apart. Type-I bursts show a characteristic spectral softening during burst decay, which was not present in any of these bursts. The spectral hardness ratios during bursts showed little variation from those of the persistent emission. A preliminary analysis of fourier power spectra shows no significant coherent or quasiperiodic oscillations between 1 and 2048 Hz in the bursts or persistent emission. Upper limits (95-percent confidence) on the strength of any narrow features in the power spectra are 3.4-percent rms. This active phase followed the previous one by 7 months (IAUC 6506). Infrared and radio observations are encouraged, particularly if coordinated with the ongoing public RXTE PCA observations.”

1997 June 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green